• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Kelleher

    Log entry time 08:11:18 on July25,2005

    Entry number 149399

    keyword=Owl Shift Summary

    00:00 Started shift in middle of HRS-P0 scan

    Time Left HRS Right HRS Run#
    00:11 2.734 2.730 2261
    00:25 2.736 2.732 2262
    00:47 2.738 2.734 2263
    00:54 2.734 2.730 2264
    01:08 2.736 2.732 2265
    01:38 2.738 2.734 2266
    01:49 2.734 2.730 2267
    02:10 2.740 2.736 2268, 2269
    03:42 2.736 2.732 2270
    03:58 2.737 2.733

    Septum Currents
    Time Left Right
    00:25 283.57 278.22
    01:08 309.35 303.52
    03:41 283.57 278.22

    02:15 MCC called: Beam down for a few minutes due to problems with the
    RF zone. Ended up being down for about 60 minutes.

    04:30 Convert DAQ back to HAPPEX

    04:44 Asked for 5uA beam to check raster.

    04:49 Spot++ checked out, asked for 35uA

    04:49 Started taking production data

    05:04 MCC wanted to send tune beam, so we moved the target to empty.

    06:50 Took HAPPEX production data

    07:24 After running at 35uA for 1/2 hour, we increased the beam current to 40uA to watch for septum heating.

    07:48 Right septum magnet quenched (at about 7.7 . Asked MCC to take beam away.

    Happex Run Summary

    Run# Events
    3812 27.6k
    3813 50.4k (adjusted high voltage)
    3814 61.2k
    3815 45.5k